find the slope and y intercept of the line and graph
x + y = 0
Thanks

Answer by jainenderkapoor(61)

To find the slope of an equation - we need to convert the equation in the slope-intercept form
y = m*x + c and compare values of `m` and `c`
so that we can get the value of slope and the `y` intercept
`m` is the slope of equation and `c` is the y-intercept
So converting the given equation in the slope-intercept form we get
y = -x + 0
y = -1*x + 0
On comparing we get values of `m` = -1 and `c` = 0
That means the line has a slope -1 and the y-intercept is 0
HOPE it is clear to you. In case you have any other doubt you can contact me on kapoorjai1@rediffmail.com
I am available for Online Tutoring Also.